OpenStudy (zepp):

Explain...?\[\large \frac{\cos \theta -1}{\theta}=-\frac{1- \cos \theta}{\theta}\]

OpenStudy (zepp):

Is it like \(-(1-\cos \theta)=-1+\cos \theta\)

OpenStudy (zepp):

Got it, thanks guys.

OpenStudy (anonymous):

Yeah...then use: \[\frac{-a}{b}=-\frac{a}{b}\]